OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3x − y = 7? Checking Answers y = 7x − 3 y = −3x + 7 -->y = −7x + 3 y = 3x

OpenStudy (ranga):

First step is to put the equation of the line 3x - y = 7 in the slope-intercept form. Slope intercept form is: y = mx + b

OpenStudy (anonymous):

Okay so y = 3x - 7

OpenStudy (ranga):

Yes. Now what is the slope of the line: y = 3x - 7 ? Remember, in y = mx + b, m is the slope and b is the y-intercept.

OpenStudy (ranga):

Compare y = 3x - 7 to y = mx + b What is the value of m? (m is called the slope)

OpenStudy (anonymous):

7/3 is the slope?

OpenStudy (ranga):

No, if you compare y = 3x - 7 to y = mx + b you can see m = 3 and b = -7 So the slope of the line is 3.

OpenStudy (anonymous):

Okay that makes sense

OpenStudy (ranga):

Two lines will be parallel if they have the same slope. Which of the 4 choices have a slope of 3?

OpenStudy (anonymous):

2 and 4 ?

OpenStudy (ranga):

2 has a slope of -3 NOT +3

OpenStudy (anonymous):

So 1 and 4 then?

OpenStudy (anonymous):

im going to go with one thanks for helping

OpenStudy (ranga):

No. The answer is 4! Earlier you narrowed it down to 2 & 4 and I said it is not 2. So that leaves you with 4!

OpenStudy (anonymous):

Okay sorry , I just double checked and it didn't come out right so I came back thanks